OpenStudy (anonymous):

find tg\he value of c that makes each trinomial a perfect square. then write the trinomial as a perfect square. the equation , x^2+12x+c

OpenStudy (anonymous):

(x+a)^2= x^2+2ax+a^2 therefore 2a=12 and c=a^2

OpenStudy (anonymous):

u got it zel?

OpenStudy (anonymous):

half of 12 is 6, 6 squared is 36 that is what you want for c

OpenStudy (anonymous):

yes.